MICROWAVE SPECTRA OF THE Ar_{3}-NH_{3} AND Ne_{3}-NH_{3} VAN DER WAALS TETRAMERS

\underline{JENNIFER VAN WIJNGAARDEN} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of the Ar_{3}-NH_{3} and Ne_{3}-NH_{3} van der Waals complexes were measured between 4 and 18 GHz using a pulsed jet Fourier transform microwave spectrometer. The isotopomers studied include those of NH_{3}, ^{15}NH_{3}, ND_{3}, ND_{2}H, and NDH_{2} in combination with Ar_{3}, ^{20}Ne_{3}, ^{22}Ne_{3}, ^{20}Ne_{2}^{22}Ne, and ^{20}Ne^{22}Ne_{2}. The isotopomers containing Ar_{3}, 20Ne_{3}, and ^{22}Ne_{3} are symmetric tops for which a-type transitions with K=3n (n=0,1,2,...) were observed. The other isotopomers are asymmetric tops and due to the reduced symmetry, all rotational levels are present in these complexes. For ^{20}Ne_{2}^{22}Ne, both a- and b-type transitions were observed whereas the slightly different mass distribution in ^{20}Ne_{2}^{22}Ne led to the detection of a- and c- type transitions. For the deuterium containing isotopomers, each rotational transition is split by the inversion motion of ammonia within the complex. This inverison splitting is not observed in the NH_{3} and 15NH_{3} containing species for spin statistical reasons. The spectroscopic constants, including the ^{14}N nuclear quadrupole coupling constants, were fit for each isotopomer and used to derive structural and dynamical information for the Ar_{3}-NH_{3} and Ne_{3}-NH_{3} van der Waals complexes.